OpenStudy (anonymous):

Find tan θ if sec θ = sqrt 10/ 3 and sin θ < 0.

OpenStudy (cwrw238):



OpenStudy (cwrw238):

y = sqrt(10 + 3^2) = sqrt19 the angle is in 4th quadrant because sec is positive and sin negative can you continue?

OpenStudy (anonymous):

yeah i did all of that, i was only confused because my answer choices are -3 3 -1/3 - sqrt 10

OpenStudy (cwrw238):

oopps sorry y = sqrt ( 10 - 3^2) = 1

OpenStudy (anonymous):

oh okay i get it now

OpenStudy (cwrw238):

-1 not 1 and tan = -1/3

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

got there eventually - welcome